1. A sample of tris(hyroxymethyl)aminomethane (H2NC(CH2OH)3), which is a base, with a mass of 0.1631 g is dissolved in water and used to standardize a solution of hydrochloric acid. An endpoint is reached when 47.60 mL of hydrochloric acid has been added. Determine the molarity (in mol/L) of the hydrochloric acid solution. Report your answer to four significant figures.
Answer #1

Gram equivalent of base =gram equivalent of acid

Mole × n-factor = M× V

0.1631/121.14×1 = M×47.60/1000

M = 0.02822 mol/L